# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ B.4 Channel-Scaled Convolutional Neural Networks
$\mathbf{\bm{\theta}}=\left(\left\{\mathbf{K}^{\left(l\right)}\right\}_{l=1}^{L},\,\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\,\left\{\bm{\gamma}^{\left(l\right)}\right\}_{l=1}^{L},\,\mathbf{w}^{\left(L+1\right)},\,b^{\left(L+1\right)}\right)\,,$
where $\mathbf{K}_{i,j}^{\left(l\right)}\in\mathbb{R}^{k_{l}}$ are convolution operators defined by kernels with $k_{l}$ parameters, $\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{c_{l}}$ are bias terms, and $\bm{\gamma}^{\left(l\right)}$ are channel scaling parameters. $\mathbf{w}^{\left(L+1\right)}\in\mathbb{R}^{s}$ and $b^{\left(L+1\right)}\in\mathbb{R}$ are the weights and bias of the convolutional network’s last fully connected layer, where $d_{s}$ is the dimension of $\mathrm{Vec}\left(\mathbf{f}_{\mathrm{SCN}}^{\left(L\right)}\left(\mathbf{x}\right)\right)$. We denote the class of all SCNs with multi-index widths $D$ as $\mathcal{H}^{\mathrm{SCN}}_{D}$.
Theorem B.12 (Teacher equivalence probability for SCN restated).
For any activation function and any depth $L$, let $D^{\star}=\left(c_{1}^{\star},\dots,c_{L}^{\star},d_{s}^{\star}\right),\;D=\left(c_{1},\dots,c_{L},d_{s}\right)\in\mathbb{N}^{L+1}$ such that $D^{\star}\leq D$. If there exists some teacher $h_{\bm{\theta}^{\star}}\in\mathcal{H}^{\mathrm{SCN}}_{D^{\star}}$ and the prior is $\mathcal{P}=\mathcal{P}\left(\mathcal{H}^{\mathrm{SCN}}_{D}\right)$ then
@@
And by Theorem 3.2, $N=(\tilde{C}+3\log 2/\delta)/\epsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\epsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
The proof is completely analogous to the proof of Theorem B.6. Specifically, we claim that
=$\mathcal{E}=\left\{h_{\bm{\theta}}\in\mathcal{H}^{\mathrm{SCN}}_{D}\middle|\begin{array}[]{l}\forall l=1,\dots,L\;\mathbf{K}^{\left(l\right),\left(11\right)}=\mathbf{K}^{\star\left(l\right)},\;\bm{\gamma}_{1}^{\left(l\right)}=\bm{\gamma}_{1}^{\star\left(l\right)},\;\bm{\gamma}_{2}^{\left(l\right)}=\mathbf{0},\\ \mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\;\mathbf{w}^{\left(L+1\right),1}=\mathbf{w}^{\star\left(L+1\right)},\;b^{\left(L+1\right)}=b^{\star\left(L+1\right)}\end{array}\right\}\,.$
@@
Appendix C Proof for the Interpolation Probability in the Continuous Case (Section 6)
C.1 Technical Lemmas
@@
Lemma C.1.
For any vector $\mathbf{y}$ and random vector $\mathbf{x}\sim\left(\mathbf{0},\mathbf{I}_{d}\right)$, $\epsilon\in\left(0,\frac{\pi}{2}\right)$ and $u\in\left(0,1\right)$ we have
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\epsilon\right)\right)=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\epsilon\right)\right)\geq\frac{\sin\left(\epsilon\right)^{d-1}}{\left(d_{0}-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}$ (41)
@@
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\epsilon\right)\right)=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\epsilon\right)\right)\,,$
and we can set $\mathbf{y}=\left[1,0\dots,0\right]^{\top}$, without loss of generality. Therefore, as in TV [2017]
@@
Therefore, for $\epsilon>0$,
$\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|^{2}>\cos^{2}\left(\epsilon\right)\right)$ $\geq\frac{2\left(1-\cos^{2}\left(\epsilon\right)\right)^{\frac{d-1}{2}}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}=\frac{2\sin\left(\epsilon\right)^{d-1}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}\,,$
@@
Corollary C.3.
For any vector $\mathbf{y}$ and $\mathbf{x}\sim\left(\mathbf{0},\mathbf{I}_{d}\right)$, $\epsilon\in\left(0,\frac{\pi}{2}\right)$ and $u\in\left(0,1\right)$ we have
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\epsilon\right)\right)$ $=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\epsilon\right)\right)$
$\geq\exp\left(d\log\left(\sin\left(\epsilon\right)\right)-\frac{1}{2}\log\left(d\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)\,.$
@@
$\frac{\sin\left(\epsilon\right)^{d-1}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}$ $=\frac{\sin\left(\epsilon\right)^{d-1}}{\left(d-1\right)\sqrt{\frac{2\pi}{d-1}}\left(1+O\left(\frac{1}{d-1}\right)\right)}$
$=\frac{\sin\left(\epsilon\right)^{d-1}}{\sqrt{2\pi\left(d-1\right)}}\left(1+O\left(d^{-1}\right)\right)$
$=\frac{\sin\left(\epsilon\right)^{d-1}}{\sqrt{2\pi\left(d-1\right)}}\left(1+O\left(d^{-1}\right)\right)$
$=\exp\left(\left(d-1\right)\log\left(\sin\left(\epsilon\right)\right)-\frac{1}{2}\log\left(d-1\right)-\frac{1}{2}d_{1}^{\star}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)$
$=\exp\left(-\left(d-1\right)\left|\log\left(\sin\left(\epsilon\right)\right)\right|-\frac{1}{2}\log\left(d-1\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)$
$>\exp\left(-d\left|\log\left(\sin\left(\epsilon\right)\right)\right|-\frac{1}{2}\log\left(d\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)\,.$
@@
C.2 Setting and Notation
Let $h_{\bm{\theta}},h_{\bm{\theta}^{\star}}$ be fully connected (Def. 5.1) two layer neural network models with input dimension $d_{0}$, output dimension $d_{2}=1$ and hidden layer dimensions $d_{1}$ and $d_{1}^{\star}$, respectively. To simplify notation, we omit the $\operatorname{sign}$ activation from the definition of $h_{\bm{\theta}}$ and $h_{\bm{\theta}^{\star}}$ and denote
$h_{\bm{\theta}}\left(\mathbf{x}\right)=\mathbf{z}^{\top}\sigma\left(\mathbf{W}\mathbf{x}\right)$
$h_{\bm{\theta}^{\star}}\left(\mathbf{x}\right)=\mathbf{z}^{\star\top}\sigma\left(\mathbf{W}^{\star}\mathbf{x}\right)$
@@ C.2 Setting and Notation
and $\sigma\left(\cdot\right)$ is the common leaky rectifier linear unit (LReLU, Maas et al. [2013]) with parameter $\rho\notin\{0,1\}$.
@@
$h_{\bm{\theta}}\left(\mathbf{X}\right)=\sigma\left(\mathbf{W}\mathbf{X}\right)^{\top}\mathbf{z}\in\mathbb{R}^{N}.$ (45)
$h_{\bm{\theta}^{\star}}\left(\mathbf{X}\right)=\sigma\left(\mathbf{W}^{\star}\mathbf{X}\right)^{\top}\mathbf{z}^{\star}\in\mathbb{R}^{N}.$ (46)
We denote the labels $y^{\left(n\right)}=\operatorname{sign}\left(h_{\bm{\theta}^{\star}}\left(\mathbf{x}^{\left(n\right)}\right)\right)$, and
@@
We use the notation $\bm{a}^{\left(n\right)}=a\left(\mathbf{W}\mathbf{x}^{\left(n\right)}\right)\in\left\{\rho,1\right\}^{d_{1}}$ for the activation pattern of the hidden layer of $h_{\bm{\theta}}$ on the input $\mathbf{x}^{\left(n\right)}$. In addition, we define the flattened weights’ vectors of $h_{\bm{\theta}}$ and $h_{\bm{\theta}^{\star}}$ as
@@
respectively, where $\mathbf{W}^{\star}$ and $\mathbf{z}^{\star}$ are padded with 00’s to match the dimensions. Let $\bm{\phi}^{\left(n\right)}=\left(\bm{a}^{\left(n\right)}\otimes\mathbf{x}^{\left(n\right)}\right)y^{\left(n\right)}\in\mathbb{R}^{d_{0}d_{1}}$ where $\otimes$ is the Kronecker product. With this notation, it can be shown that
@@
Assumption C.4 (Prior over hypotheses, continuous setting, restated).
Suppose that the weights of $h_{\bm{\theta}}$ are random such that each row of the first layer, $\mathbf{w}_{i}$, is independently sampled from a uniform distribution on the unit sphere $\mathbb{S}^{d_{0}-1}$, and the second layer $\mathbf{z}$ is sampled uniformly from $\mathbb{S}^{d_{1}-1}$. Both $\mathbf{w}_{i}$ and $\mathbf{z}$ are independent of the teacher and data.
C.3 Proof of Theorem 6.7
Definition C.5 (First layer angular margin, restated).
@@
We denote the event that $h_{\bm{\theta}^{\star}}$ and $h_{\bm{\theta}}$ agree on the activation pattern of the data by
@@
To bound the probability of this event, we use the following Lemma, adapted from Soudry and Hoffer [2017]. For completeness, we write its proof here.
Lemma C.6 (Activation matching probability, adapted from Soudry and Hoffer [2017]).
Let $\mathbf{X}$ be a dataset with first layer angular margin $\alpha$ w.r.t $h_{\bm{\theta}^{\star}}$. Then
@@
where in (1) we used (49), in (2) we used the independence of $\left\{\mathbf{w}_{i}\right\}_{i=1}^{d_{1}^{\star}}$ and in (3) we used (C.1). When $d_{0}\gg 1$, we can use Corollary C.3 to get
@@
We can simplify this equation when $d_{0}\gg 1$ with the asymptotic expansion of the beta function from Lemma C.2. If $d_{1}^{\star}d_{0}^{-1}\ll 0$ then the error $\left(1+O\left(d_{0}^{-1}\right)\right)^{d_{1}^{\star}}=1+O\left(d_{0}^{-1}d_{1}^{\star}\right)$. Overall, this means that
@@
Definition C.7 (Second layer angular margin, restated).
@@
$\forall n\in\left[N\right]\;\left|\frac{h_{\bm{\theta}^{\star}}\left(\mathbf{x}^{\left(n\right)}\right))}{\left\|\mathbf{x}^{\left(n\right)}\right\|_{2}\left\|\mathbf{z}^{\star}\right\|_{2}}\right|>\sqrt{d_{1}\left(1+\rho^{2}\right)}\sin\beta\,.$ (54)
Assumption C.8.
Let $\alpha<\beta\in\left(0,\frac{\pi}{2}\right)$. There exists $\lambda\in\left(0,1\right)$ such that, with probability at least $1-\lambda$ over the training set $\mathcal{S}=\{\left(\mathbf{x}_{n}\right)\}_{n=1}^{N}\sim\mathcal{D}^{N}$, $\mathcal{S}$ has first layer angular margin $\alpha$ (Def. C.5) and second layer angular margin $\beta$ (Def. C.7).
@@
Theorem C.9 (Interpolation of Continuous Networks, restated).
Under Assumption C.8, with probability at least $1-\lambda$ over the dataset $\mathcal{S}$, the probability of interpolation is lower-bounded by
$p_{\mathcal{S}}$ $=\mathbb{P}_{\mathbf{W},\mathbf{z}}\left(\mathcal{L}_{S}\left(h\right)=0\right)$
@@
By Assumption C.8 with probability at least $1-\lambda$, $\mathbf{X}$ has first and second layers angular margins $\alpha$ and $\beta$, respectively, w.r.t $h_{\bm{\theta}^{\star}}$. We assume that these properties hold for the rest of the proof. Suppose that we condition on the event $\bigcap_{i=1}^{d_{1}^{\star}}\tilde{\mathcal{G}}_{i}^{\alpha}\left(\mathbf{W}^{\star}\right)$ from (49). Then, as in the proof of Lemma C.6, under the margin assumptions, $h_{\bm{\theta}}$ and $h_{\bm{\theta}^{\star}}$ agree on the activation pattern $\mathbf{a}^{\left(n\right)}$ and therefore
@@
$\left\|\mathbf{\bm{\phi}}^{\left(n\right)}\right\|^{2}=\left\|\mathbf{a}^{\left(n\right)}\right\|^{2}\left\|\mathbf{x}^{\left(n\right)}\right\|^{2}\,.$
@@
Following the same logic as in the proof of Lemma C.6, in order that $\operatorname{sign}\left(y^{\left(n\right)}\right)\neq\operatorname{sign}\left(h\left(\mathbf{x}^{\left(n\right)}\right)\right)$, $\mathbf{w}$ must be rotated by angle at least $\beta$ compared to $\mathbf{w}^{\star}$. That is,
@@
From Lemma C.1, we obtain
@@
As for (59), we know from Lemma C.6 that
@@
When $d_{0}\gg d_{1}^{\star}\gg 1$ we get from Corollary C.3
@@
$p_{\mathcal{S}}$ $=\mathbb{P}_{\mathbf{w}}\left(\forall n\in\left[N\right]\,y^{\left(n\right)}h\left(\mathbf{x}^{\left(n\right)}\right)>0\right)$
@@
The following generalization bound follows directly from Prop. A.19 and Theorem 6.7.
Corollary C.10 (Generalization of continuous two layer networks, restated).
Under the assumption that $p_{\mathcal{S}}<\frac{1}{2}$, for any $\epsilon,\delta\in\left(0,1\right)$,
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\ -\lambda,,$
@@
$N\geq\frac{\hat{C}^{\mathrm{cont}}+4\log\left(\frac{8}{\delta}\right)+2\log\left(\hat{C}^{\mathrm{cont}}\right)}{\epsilon}$

### v2

@@ C.4 Channel-Scaled Convolutional Neural Networks
$\mathbf{\boldsymbol{\theta}}=\left(\left\{\mathbf{K}^{\left(l\right)}\right\}_{l=1}^{L},\,\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\,\left\{\boldsymbol{\gamma}^{\left(l\right)}\right\}_{l=1}^{L},\,\mathbf{w}^{\left(L+1\right)},\,b^{\left(L+1\right)}\right)\,,$
where $\mathbf{K}_{i,j}^{\left(l\right)}\in\mathbb{R}^{k_{l}}$ are convolution operators defined by kernels with $k_{l}$ parameters, $\mathbf{b}^{\left(l\right)}\in\mathbb{R}^{c_{l}}$ are bias terms, and $\boldsymbol{\gamma}^{\left(l\right)}$ are channel scaling parameters. $\mathbf{w}^{\left(L+1\right)}\in\mathbb{R}^{s}$ and $b^{\left(L+1\right)}\in\mathbb{R}$ are the weights and bias of the convolutional network’s last fully connected layer, where $d_{s}$ is the dimension of $\mathrm{Vec}\left(\mathbf{f}_{\mathrm{SCN}}^{\left(L\right)}\left(\mathbf{x}\right)\right)$. We denote the class of all SCNs with multi-index widths $D$ as $\mathcal{H}^{\mathrm{SCN}}_{D}$.
Theorem C.12 (Teacher equivalence probability for SCN restated).
For any activation function and any depth $L$, let $D^{\star}=\left(c_{1}^{\star},\dots,c_{L}^{\star},d_{s}^{\star}\right),\;D=\left(c_{1},\dots,c_{L},d_{s}\right)\in\mathbb{N}^{L+1}$ such that $D^{\star}\leq D$. If there exists some teacher $h_{\boldsymbol{\theta}^{\star}}\in\mathcal{H}^{\mathrm{SCN}}_{D^{\star}}$ and the prior is $\mathcal{P}=\mathcal{P}\left(\mathcal{H}^{\mathrm{SCN}}_{D}\right)$ then
@@
And by Lemma 3.2, $N=(\tilde{C}+3\log 2/\delta)/\varepsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\varepsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
The proof is completely analogous to the proof of Theorem C.6. Specifically, we claim that
=$\mathcal{E}=\left\{h_{\boldsymbol{\theta}}\in\mathcal{H}^{\mathrm{SCN}}_{D}\middle|\begin{array}[]{l}\forall l=1,\dots,L\;\mathbf{K}^{\left(l\right),\left(11\right)}=\mathbf{K}^{\star\left(l\right)},\;\boldsymbol{\gamma}_{1}^{\left(l\right)}=\boldsymbol{\gamma}_{1}^{\star\left(l\right)},\;\boldsymbol{\gamma}_{2}^{\left(l\right)}=\mathbf{0},\\ \mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\;\mathbf{w}^{\left(L+1\right),1}=\mathbf{w}^{\star\left(L+1\right)},\;b^{\left(L+1\right)}=b^{\star\left(L+1\right)}\end{array}\right\}\,.$
@@
Appendix D Proof for the Interpolation Probability in the Continuous Case (Section 5)
D.1 Technical Lemmas
@@
Lemma D.1.
For any vector $\mathbf{y}$ and random vector $\mathbf{x}\sim\left(\mathbf{0},\mathbf{I}_{d}\right)$, $\varepsilon\in\left(0,\frac{\pi}{2}\right)$ and $u\in\left(0,1\right)$ we have
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\varepsilon\right)\right)=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\varepsilon\right)\right)\geq\frac{\sin\left(\varepsilon\right)^{d-1}}{\left(d_{0}-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}$ (42)
@@
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\varepsilon\right)\right)=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\varepsilon\right)\right)\,,$
and we can set $\mathbf{y}=\left[1,0\dots,0\right]^{\top}$, without loss of generality. Therefore, as in (TV, 2017)
@@
Therefore, for $\varepsilon>0$,
$\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|^{2}>\cos^{2}\left(\varepsilon\right)\right)$ $\geq\frac{2\left(1-\cos^{2}\left(\varepsilon\right)\right)^{\frac{d-1}{2}}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}=\frac{2\sin\left(\varepsilon\right)^{d-1}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}\,,$
@@
Corollary D.3.
For any vector $\mathbf{y}$ and $\mathbf{x}\sim\left(\mathbf{0},\mathbf{I}_{d}\right)$, $\varepsilon\in\left(0,\frac{\pi}{2}\right)$ and $u\in\left(0,1\right)$ we have
$\mathbb{P}\left(\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}>\cos\left(\varepsilon\right)\right)$ $=\frac{1}{2}\mathbb{P}\left(\left|\frac{\mathbf{x}^{\top}\mathbf{y}}{\left\|\mathbf{x}\right\|\left\|\mathbf{y}\right\|}\right|>\cos\left(\varepsilon\right)\right)$
$\geq\exp\left(d\log\left(\sin\left(\varepsilon\right)\right)-\frac{1}{2}\log\left(d\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)\,.$
@@
$\frac{\sin\left(\varepsilon\right)^{d-1}}{\left(d-1\right)B\left(\frac{1}{2},\frac{d-1}{2}\right)}$ $=\frac{\sin\left(\varepsilon\right)^{d-1}}{\left(d-1\right)\sqrt{\frac{2\pi}{d-1}}\left(1+O\left(\frac{1}{d-1}\right)\right)}$
$=\frac{\sin\left(\varepsilon\right)^{d-1}}{\sqrt{2\pi\left(d-1\right)}}\left(1+O\left(d^{-1}\right)\right)$
$=\frac{\sin\left(\varepsilon\right)^{d-1}}{\sqrt{2\pi\left(d-1\right)}}\left(1+O\left(d^{-1}\right)\right)$
$=\exp\left(\left(d-1\right)\log\left(\sin\left(\varepsilon\right)\right)-\frac{1}{2}\log\left(d-1\right)-\frac{1}{2}d_{1}^{\star}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)$
$=\exp\left(-\left(d-1\right)\left|\log\left(\sin\left(\varepsilon\right)\right)\right|-\frac{1}{2}\log\left(d-1\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)$
$>\exp\left(-d\left|\log\left(\sin\left(\varepsilon\right)\right)\right|-\frac{1}{2}\log\left(d\right)-\frac{1}{2}\log\left(2\pi\right)\right)\left(1+O\left(d^{-1}\right)\right)\,.$
@@
D.2 Setting and Notation
Let $h_{\boldsymbol{\theta}},h_{\boldsymbol{\theta}^{\star}}$ be fully connected (Def. 4.1) two layer neural network models with input dimension $d_{0}$, output dimension $d_{2}=1$ and hidden layer dimensions $d_{1}$ and $d_{1}^{\star}$, respectively. To simplify notation, we omit the $\operatorname{sign}$ activation from the definition of $h_{\boldsymbol{\theta}}$ and $h_{\boldsymbol{\theta}^{\star}}$ and denote
$h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=\mathbf{z}^{\top}\sigma\left(\mathbf{W}\mathbf{x}\right)$
$h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}\right)=\mathbf{z}^{\star\top}\sigma\left(\mathbf{W}^{\star}\mathbf{x}\right)$
@@ D.2 Setting and Notation
and $\sigma\left(\cdot\right)$ is the common leaky rectifier linear unit (LReLU, Maas et al. (2013)) with parameter $\rho\notin\{0,1\}$.
@@
$h_{\boldsymbol{\theta}}\left(\mathbf{X}\right)=\sigma\left(\mathbf{W}\mathbf{X}\right)^{\top}\mathbf{z}\in\mathbb{R}^{N}.$ (46)
$h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{X}\right)=\sigma\left(\mathbf{W}^{\star}\mathbf{X}\right)^{\top}\mathbf{z}^{\star}\in\mathbb{R}^{N}.$ (47)
We denote the labels $y^{\left(n\right)}=\operatorname{sign}\left(h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}^{\left(n\right)}\right)\right)$, and
@@
We use the notation $\boldsymbol{a}^{\left(n\right)}=a\left(\mathbf{W}\mathbf{x}^{\left(n\right)}\right)\in\left\{\rho,1\right\}^{d_{1}}$ for the activation pattern of the hidden layer of $h_{\boldsymbol{\theta}}$ on the input $\mathbf{x}^{\left(n\right)}$. In addition, we define the flattened weights’ vectors of $h_{\boldsymbol{\theta}}$ and $h_{\boldsymbol{\theta}^{\star}}$ as
@@
respectively, where $\mathbf{W}^{\star}$ and $\mathbf{z}^{\star}$ are padded with 00’s to match the dimensions. Let $\boldsymbol{\phi}^{\left(n\right)}=\left(\boldsymbol{a}^{\left(n\right)}\otimes\mathbf{x}^{\left(n\right)}\right)y^{\left(n\right)}\in\mathbb{R}^{d_{0}d_{1}}$ where $\otimes$ is the Kronecker product. With this notation, it can be shown that
@@
Assumption D.4 (Prior over hypotheses, continuous setting, restated).
Suppose that the weights of $h_{\boldsymbol{\theta}}$ are random such that each row of the first layer, $\mathbf{w}_{i}$, is independently sampled from a uniform distribution on the unit sphere $\mathbb{S}^{d_{0}-1}$, and the second layer $\mathbf{z}$ is sampled uniformly from $\mathbb{S}^{d_{1}-1}$. Both $\mathbf{w}_{i}$ and $\mathbf{z}$ are independent of the teacher and data.
D.3 Proof of Theorem 5.6
Definition D.5 (First layer angular margin, restated).
@@
We denote the event that $h_{\boldsymbol{\theta}^{\star}}$ and $h_{\boldsymbol{\theta}}$ agree on the activation pattern of the data by
@@
To bound the probability of this event, we use the following Lemma, adapted from Soudry & Hoffer (2017). For completeness, we write its proof here.
Lemma D.6 (Activation matching probability, adapted from Soudry & Hoffer (2017)).
Let $\mathbf{X}$ be a dataset with first layer angular margin $\alpha$ w.r.t $h_{\boldsymbol{\theta}^{\star}}$. Then
@@
where in (1) we used (50), in (2) we used the independence of $\left\{\mathbf{w}_{i}\right\}_{i=1}^{d_{1}^{\star}}$ and in (3) we used (D.1). When $d_{0}\gg 1$, we can use Corollary D.3 to get
@@
We can simplify this equation when $d_{0}\gg 1$ with the asymptotic expansion of the beta function from Lemma D.2. If $d_{1}^{\star}d_{0}^{-1}\ll 0$ then the error $\left(1+O\left(d_{0}^{-1}\right)\right)^{d_{1}^{\star}}=1+O\left(d_{0}^{-1}d_{1}^{\star}\right)$. Overall, this means that
@@
Definition D.7 (Second layer angular margin, restated).
@@
$\forall n\in\left[N\right]\;\left|\frac{h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}^{\left(n\right)}\right))}{\left\|\mathbf{x}^{\left(n\right)}\right\|_{2}\left\|\mathbf{z}^{\star}\right\|_{2}}\right|>\sqrt{d_{1}\left(1+\rho^{2}\right)}\sin\beta\,.$ (55)
Assumption D.8.
Let $\alpha<\beta\in\left(0,\frac{\pi}{2}\right)$. There exists $\lambda\in\left(0,1\right)$ such that, with probability at least $1-\lambda$ over the training set $\mathcal{S}=\{\left(\mathbf{x}_{n}\right)\}_{n=1}^{N}\sim\mathcal{D}^{N}$, $\mathcal{S}$ has first layer angular margin $\alpha$ (Def. D.5) and second layer angular margin $\beta$ (Def. D.7).
@@
Theorem D.9 (Interpolation of Continuous Networks, restated).
Under Assumption D.8, with probability at least $1-\lambda$ over the dataset $\mathcal{S}$, the probability of interpolation is lower-bounded by
$\hat{p}_{\mathcal{S}}$ $=\mathbb{P}_{\mathbf{W},\mathbf{z}}\left(\mathcal{L}_{S}\left(h\right)=0\right)$
@@
By Assumption D.8 with probability at least $1-\lambda$, $\mathbf{X}$ has first and second layers angular margins $\alpha$ and $\beta$, respectively, w.r.t $h_{\boldsymbol{\theta}^{\star}}$. We assume that these properties hold for the rest of the proof. Suppose that we condition on the event $\bigcap_{i=1}^{d_{1}^{\star}}\tilde{\mathcal{G}}_{i}^{\alpha}\left(\mathbf{W}^{\star}\right)$ from (50). Then, as in the proof of Lemma D.6, under the margin assumptions, $h_{\boldsymbol{\theta}}$ and $h_{\boldsymbol{\theta}^{\star}}$ agree on the activation pattern $\mathbf{a}^{\left(n\right)}$ and therefore
@@
$\left\|\mathbf{\boldsymbol{\phi}}^{\left(n\right)}\right\|^{2}=\left\|\mathbf{a}^{\left(n\right)}\right\|^{2}\left\|\mathbf{x}^{\left(n\right)}\right\|^{2}\,.$
@@
Following the same logic as in the proof of Lemma D.6, in order that $\operatorname{sign}\left(y^{\left(n\right)}\right)\neq\operatorname{sign}\left(h\left(\mathbf{x}^{\left(n\right)}\right)\right)$, $\mathbf{w}$ must be rotated by angle at least $\beta$ compared to $\mathbf{w}^{\star}$. That is,
@@
From Lemma D.1, we obtain
@@
As for (60), we know from Lemma D.6 that
@@
When $d_{0}\gg d_{1}^{\star}\gg 1$ we get from Corollary D.3
@@
$\hat{p}_{\mathcal{S}}$ $=\mathbb{P}_{\mathbf{w}}\left(\forall n\in\left[N\right]\,y^{\left(n\right)}h\left(\mathbf{x}^{\left(n\right)}\right)>0\right)$
@@
The following generalization bound follows directly from Prop. B.23 and Theorem 5.6.
Corollary D.10 (Generalization of continuous two layer networks, restated).
Under the assumption that $\hat{p}_{\mathcal{S}}<\frac{1}{2}$, for any $\varepsilon,\delta\in\left(0,1\right)$,
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\ -\lambda,,$
@@
$N\geq\frac{\hat{C}^{\mathrm{cont}}+4\log\left(\frac{8}{\delta}\right)+2\log\left(\hat{C}^{\mathrm{cont}}\right)}{\varepsilon}$